OpenStudy (compassionate):

Why is "Simplify" not an actual mathematical term? Something so intuitive must have some core terminology to it.

OpenStudy (anonymous):

it is the lazy math teacher's way of saying "write the answer i want" sometimes it means compute the number sometimes it means combine like terms sometimes it means reduce the fraction to lowest terms sometimes it means write in "simplest radical form" which does have the word "simplest" in it, but has well defined rules there is such a thing as "canonical form" but that terminology is rarely used

OpenStudy (compassionate):

Why did I hear the term "simplify," all too often throughout my High School and Middle School times? By simplify, surely they mean reduce until it cannot be reduced further, or simplest form. e.g., 2x + x + y = 3x + y

OpenStudy (anonymous):

oh and btw there really is no such operation as "cross multiply" it is true that if \(\frac{a}{b}=\frac{c}{d}\) then \(ad=bd\) but you do not "cross multiply" when adding, subtracting etc

OpenStudy (anonymous):

what you did above is "combine like terms" is that the same as writing \(\frac{12}{8}\) as \(\frac{3}{2}\)?

OpenStudy (compassionate):

Ah, I was pulling your leg with that one, @satellite73 , @e.mccormick and I had a conversation about 6 months ago and he mentioned you hated some terms, especially cross multiply, so I decided to test that theory. Him and I are, to say, good friends.

jimthompson5910 (jim_thompson5910):

as this page puts it `"Simplified" is sometimes obvious, but can also depend on what you want to do.` basically it depends on the context and what the teacher wants. It's sometimes a subjective thing where not all agree on what the fully simplified form is https://www.mathsisfun.com/algebra/simplify.html

OpenStudy (compassionate):

Then, there is no true reason to all this terminology. It's just useless theory and logic in applied fields. We all have different ways we can ,"simplify," expressions, equations, fractions, yada-yada. It's really nonsensical given the measurement of degree the terms ,"simplify," and, "cross-multiply," are used in high school and primary arithmetic. @jim_thompson5910 , when math becomes subjective and unreasonable, then Hell freezes over.